OpenStudy (anonymous):

What is the solution of each proportion? 5/9=m/54

OpenStudy (kenljw):

5/9=m/54 54 (5/9) = m 6(5)=30=m

OpenStudy (anonymous):

what's the answer? I'm confused. @KenLJW

OpenStudy (kenljw):

When solving a equation what ever you do to one side you must do to the other to maintain equality m/54 = 5/9 54 (m/54) =(5/9) 54 m = 5 (6) = 30

OpenStudy (anonymous):

is the answer 30?

OpenStudy (anonymous):

@KenLJW

OpenStudy (kenljw):

yes